40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15602     Entry time: Wed Sep 23 15:06:54 2020
Author: Jordan 
Type: Update 
Category: VAC 
Subject: TP2 Forepump Re-install 

I removed the forepump to TP2 this morning after the vacuum failure, and tested in the C&B lab. I pumped down on a small volume 10 times, with no issue. The ultimate pressure was ~30 mtorr.

I re-installed the forepump in the afternoon, and restarted TP2, leaving V4 closed. This will run overnight to test, while TP3 backs TP1.

In order to open V1, with TP3 backing TP1, the interlock system had to be reset since it is expecting TP2 as a backing pump. TP2 is running normally, and pumping of the main volume has resumed.


gautam 2030:

  1. The monitor (LCD display) at the vacuum rack doesn't work - this has been the case since Monday at least. I usually use my laptop to ssh in so I didn't notice it so it could have been busted from before. But for anyone wishing to use the workstation arrangement at 1X8, this is not great. Today, we borrowed the vertex laptop to ssh in, the vertex laptop has since been returned to its nominal location.
  2. The modification to the interlock condition was made by simply commenting out the line requiring V4 to be open for V1 to be opened. I made a copy of the original .yaml file which we can revert to once we go back to the normal config.
  3. I also opened VM1 to allow the RGA scans to continue to be meaningful.
  4. At the time of writing, all systems seem nominal. See Attachment #2. The vertical line indicates when we started pumping on the main volume again earlier today, with TP3 backing TP1.

Unclear why the TP2 foreline pump failed in the first place, it has been running fine for several hours now (although TP2 has no load, since V4 isolates it from the main volume). Koji's plots show that the TP2 foreline pressure did not recover even after the interlock tripped and V4 was closed (i.e. the same conditions as TP2 sees right now).

Attachment 1: Screenshot_from_2020-09-23_15-15-43.png  32 kB  Uploaded Wed Sep 23 16:16:29 2020  | Hide | Hide all
Screenshot_from_2020-09-23_15-15-43.png
Attachment 2: MainVolPumpDown.png  185 kB  Uploaded Wed Sep 23 21:37:44 2020  | Hide | Hide all
MainVolPumpDown.png
ELOG V3.1.3-